(5x-5)(6x+3)=0
solve

Answer by checkley77(12844)

(5x-5)(6x+3)=0
5X-5=0
5X=5
X=5/5
X=1 ANS.
6X+3=0
6X=-3
X=-3/6
X=-1/2 ANS.